Paul Sabin organized Sabin Electronics 10 years ago to produce and sell several electronic devices on which he had secured paSabin Electronics Comparative Income Statement and Reconciliation This Year Last Year Sales $ 5,450,000 $ 4,620,000 Cost of gRequired: 1. You decide first to assess the companys stock market performance. For both this year and last year, compute: a.Complete this question by entering your answers in the tabs below. Required 1 Required 2 You decide first to assess the compaComplete this question by entering your answers in the tabs below. Required 1 Required 2 ---------- You decide next to assess
Answer #1

Earning per share = Net Income / Common Stock outstanding
Last Year = $301000/52000 = $5.79 per share
This Year = $511000/52000 = $9.83 per share
Common Stock outstanding = $780000/15 = 52000

Dividend Yield Ratio = Dividend per share / Market price
Last year = $1.88 / $55 = 3.43%
Dividend per share = $98000/52000 = $1.88
This year = $2.29 / $60 = 3.81%
Dividend per share = $119000/52000 = $2.29

Dividend payout ratio = Dividend paid / Net Income
Last year = $98000/ 301000 = 32.56%
This year = $119000/ 511000 = 23.29%

Price earning ratio = Market price / EPS
Last year = $55 / 5.79 = 9.50
This year = $60 / 9.83 = 6.10

Book Value per share = Total Stockholders Equity / Common stock outstanding
Last year = $1573000 / 52000 = $30.25 per share
This year = $1965000 / 52000 = $37.79 per share

As per HOMEWORKLIB RULES we are suppose to answer 4 parts, i have answered 5, so kindly post others separately